# Existence of Friedrich-Wintgen Bound States in the Continuum: Cavity with a Thin Waveguide Opening

### v1

@@ Abstract
Bound states in the continuum (BICs) are localized states embedded within a continuum of propagating waves. Perturbations that disrupt BICs typically induce ultra-strong resonances, a phenomenon enabling diverse applications in photonics. This work investigates the existence of BICs in two-dimensional electromagnetic cavities coupled to thin waveguides for H-polarized waves. Our focus is on Friedrich-Wintgen BICs (FW-BICs), which arise from destructive interference between two resonant modes and were identified numerically in rectangular cavities with waveguide openings by Lyapina et al. [J. Fluid Mech., 780 (2015), pp. 370–387]. Here, we rigorously establish the existence of FW-BICs in a broader class of cavity geometries by introducing perturbations to the refractive index under regularity constraints. We show that BICs correspond to intersections of two curves derived implicitly from the governing equations constructed via the mode-matching method. Crucially, we prove that such intersections are guaranteed for sufficiently small waveguide widths, provided that two eigenvalues of the cavity cross and the associated eigenfunctions exhibit non-vanishing coupling to the radiation channel at the cavity-waveguide interface. Furthermore, our approach remains applicable for studying the emergence of FW-BICs under parameter-dependent boundary perturbations to the cavity.
@@ 1 Introduction
In this work, we investigate the existence of BICs in a broader class of 2D cavities coupled to semi-infinite waveguides. We consider cavities featuring a line segment on the boundary from which a semi-infinite waveguide extends. A perturbation to the refractive index, analytically dependent on a parameter $\delta$ under specific regularity conditions, is introduced. By reducing the infinite-dimensional linear system derived from the mode-matching method to two governing equations, we show that BICs are characterized as intersections of two curves in the $(\delta,\mu)$-parameter plane, where $\mu$ denotes the square of the free-space wavenumber. Assuming the waveguide width is sufficiently small, we show these intersections are guaranteed to exist, provided that two eigenvalues of the cavity cross as $\delta$ varies and the associated eigenfunctions exhibit non-vanishing coupling to the radiation channel at the cavity-waveguide interface. The main result is formalized in Theorem 5.13. Our method remains applicable for showing the existence of BICs under parameter-dependent boundary perturbations, if the eigenvalues and eigenfunctions exhibit continuous dependence on these perturbations. This continuity condition is satisfied in specific scenarios (cf. Section VII.6.5 in [10]), including scaling transformations of rectangular cavities as explored in [16]. For additional applications of analogous mode-matching techniques in resonance analysis of subwavelength structures, we refer the readers to [14, 15, 33].
This paper is organized as follows. Section 2 outlines the problem setup and three key conditions required for the analysis. In Section 3, we construct the governing infinite-dimensional linear system. Section 4 reformulates the system into a purely real-valued form and reduces it to finite dimensions. Section 5 rigorously establishes the existence of FW-BICs. Numerical experiments validating the theoretical findings are presented in Section 6. Finally, Section 7 briefly discusses the existence of FW-BICs under boundary perturbations, alongside other potential extensions of our approach.
@@ 2 Problem formulation and key requirements
As illustrated in Fig. 1, our problem region $\Omega$ is defined as $\Omega:=\Omega_{\mathrm{in}}\cup\Gamma_{h}\cup\Omega_{\mathrm{out}}$, where $\Omega_{\mathrm{in}}$ represents the cavity region with a line segment on its boundary, $\Omega_{\mathrm{out}}$ denotes the waveguide region of width $h$ extended from the line segment, and $\Gamma_{h}$ is the common boundary between these two regions. Letting $\mathbf{o}$ denote the middle point of $\Gamma_{h}$, we establish a rectangular coordinate system in $\mathbb{R}^{2}$ centered at $\mathbf{o}$. Without loss of generality, we always fix the point $\mathbf{o}$ on $\partial\Omega_{\mathrm{in}}$ as the middle point of $\Gamma_{h}$ when $h$ is allowed to decrease in the subsequent sections. Additionally, $\Omega_{B}$ denotes a rectangular region in $\Omega_{\mathrm{in}}$ that $\Gamma_{h}\subset\partial\Omega_{B}$.
$\Omega_{\mathrm{in}}$$n(\mathbf{x})$$\Omega_{\mathrm{out}}$$h$$\Omega_{\mathrm{in}}$$n(\mathbf{x})$$\Omega_{\mathrm{out}}$$h$$\mathbf{o}$$x_{1}$$x_{2}$${\Gamma_{h}}$${\Omega_{B}}$
@@
(1) {$&n_{0}\leq{n}(\mathbf{x})\leq{n_{1}},\ &&\text{if}\ \mathbf{x}\in\Omega_{\mathrm{in}};\\ &n(\mathbf{x})=1,\ &&\text{if}\ \mathbf{x}\in\Omega_{\mathrm{out}}\cup\Omega_{B},\\ \end{aligned}\right.$
@@
(3) $-\mathcal{P}{u}+\mu{u}$ $=0\ \mathrm{in}\ \Omega,$
@@
(5) $-\mathcal{P}{\psi}+\lambda{\psi}$ $=0\ \mathrm{in}\ \Omega_{\mathrm{in}},$
@@
    At $\delta=0$, the eigenvalue functions $\lambda_{M-2}(\delta)$ and $\lambda_{M-1}(\delta)$ cross, resulting in an eigenvalue of multiplicity 2.
@@ 3.1 Eigenfunction expansion in the cavity
Under Condition A.1, we can choose a set of normalized eigenfunctions $\{\psi_{m}(\cdot,\delta)\}$ that the mapping $\delta\to\psi_{m}(\cdot,\delta)$ is continuous from $[-\delta_{0},\delta_{0}]$ into $H^{1}(\Omega_{\mathrm{in}})$ for each $m\in\mathbb{N}$, i.e.,
(8) $\|\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2})\|_{H^{1}(\Omega_{\mathrm{in}})}\to 0\ \text{as}\ \delta_{1}\to\delta_{2}\ \text{for}\ \delta_{1},\delta_{2}\in[-\delta_{0},\delta_{0}].$
@@
(9) $\left|\frac{1}{\lambda_{m}(\delta_{1})+1}-\frac{1}{\lambda_{m}(\delta_{2})+1}\right|\leq{e^{C_{0}|\delta_{1}-\delta_{2}|}-1},\ \text{for}\ \delta_{1},\delta_{2}\in[-\delta_{0},\delta_{0}],$
@@
Since $n^{2}(\cdot,\delta)$ varies analytically with respect to $\delta$ and $H^{1}(\Omega_{\mathrm{in}})$ is dense in $L^{2}(\Omega_{\mathrm{in}})$, $\Phi_{\delta}$ is a self-adjoint holomorphic family of forms of type (a) (cf. Sections VII.4.2 and VII.4.8 in [10]). Hence, $\psi_{m}(\cdot,\delta)$ can be chosen as a continuous mapping from $[-\delta_{0},\delta_{0}]$ to $L^{2}(\Omega_{\mathrm{in}})$, and its eigenvalue $\lambda_{m}(\delta)$ satisfies (9) by using the results in Section VII.4.7 of [10]. To prove (8), for each $m\in\mathbb{N}$, direct computation gives:
$\Phi_{\delta_{2}}(\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2}),\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2}))$
$=$ $(\lambda_{m}(\delta_{1})\psi_{m}(\cdot,\delta_{1})-\lambda_{m}(\delta_{2})\psi_{m}(\cdot,\delta_{2}),\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2}))_{\Omega_{\mathrm{in}}}$
(12) $+((\frac{1}{n^{2}(\cdot,\delta_{2})}-\frac{1}{n^{2}(\cdot,\delta_{1})})\nabla\psi_{m}(\cdot,\delta_{1}),\nabla(\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2})))_{\Omega_{\mathrm{in}}}.$
where the divergence theorem and the boundary conditions $\partial_{\bm{\nu}}\psi_{m}(\cdot,\delta)=0$ on $\partial\Omega_{\mathrm{in}}$ for every $\delta$ have been applied. The vanishing of (12) as $\delta_{1}\to\delta_{2}$ follows from the $L^{2}$-convergence $\|\psi_{m}(\cdot,\delta_{1})-\psi_{m}(\cdot,\delta_{2})\|_{L^{2}(\Omega_{\mathrm{in}})}\to 0$, the analytic dependence of $n^{2}(\cdot,\delta_{1})$ on $\delta_{1}$, and the uniform boundedness of $\|\psi_{m}(\cdot,\delta_{1})\|_{H^{1}(\Omega_{\mathrm{in}})}$.
@@
(17) $\|u\|_{H^{1}(\Omega_{\mathrm{in}})}=\sqrt{\|u\|^{2}_{L^{2}(\Omega_{\mathrm{in}})}+\|\frac{1}{n}\nabla{u}\|^{2}_{L^{2}(\Omega_{\mathrm{in}})}},\ \text{for}\ u{\in}H^{1}(\Omega_{\mathrm{in}}).$
@@ 3.2 Mode-expansion in the waveguide
and $u$ is supposed to be outgoing. This gives the expansions:
(20) $u=\sum_{j=0}^{\infty}b_{j}e^{{\bf i}\alpha_{j}x_{1}}\phi_{j}(x_{2}),$
@@
$\alpha_{j}=\sqrt{\mu-\left(\frac{j\pi}{h}\right)^{2}}$. We also introduce a set of functions $\{\eta_{j}\}$
(22) $\eta_{j}:=\sqrt{\frac{2}{h}}\sin\left(\frac{j\pi}{h}x_{2}+\frac{j\pi}{2}\right),\ j\in\mathbb{Z}_{+}.$
@@
(23) $u|_{\Gamma_{h}}=$ $\sum_{j=0}^{\infty}b_{j}\phi_{j}$ $\text{in}\ H^{1/2}(\Gamma_{h}),$
(24) $\partial_{\bm{\nu}}u|_{\Gamma_{h}}=$ $\sum_{j=0}^{\infty}b_{j}{\bf i}\alpha_{j}\phi_{j}$ $\text{in}\ \widetilde{H}^{-1/2}(\Gamma_{h}).$
@@ 4 Reduction of the system
Given an integer $M\geq 3$, we make the following assumptions when $\delta=0$:
(30) $\lambda_{M-3}(0)<\lambda_{M-2}(0)=\lambda_{M-1}(0)<\lambda_{M}(0).$
Then the following properties regarding the stability of $\{\lambda_{m}(\delta)\}$ hold.
@@
Under Condition A.1, for every $\epsilon>0$, there exists a closed interval for $\delta$ such that
@@
By selecting a range for $\delta$ such that
(35) $e^{C_{0}\delta}-1\leq\frac{1}{\lambda_{M}(0)+1-\epsilon}-\frac{1}{\lambda_{M}(0)+1},$
@@
(36) $\epsilon<\frac{1}{3}\min\Big{\{}\lambda_{M-2}(0)-\lambda_{M-3}(0),\lambda_{M}(0)-\lambda_{M-1}(0)\Big{\}}.$
In the following, we assume $\delta$ varies within a range where (31)–(32) hold and
@@
Our goal is to reformulate the system (28) into a finite-dimensional real-valued system for $h\ll 1$.

### v2

@@ Abstract
Bound states in the continuum (BICs) are localized states embedded within a continuum of propagating waves. Perturbations that disrupt BICs typically induce ultra-strong resonances, a phenomenon enabling diverse applications in photonics. This work investigates the existence of BICs in two-dimensional electromagnetic cavities coupled to thin waveguides for H-polarized waves. Our focus is on Friedrich-Wintgen BICs (FW-BICs), which arise from destructive interference between two resonant modes and were identified numerically in rectangular cavities with waveguide openings by Lyapina et al. [J. Fluid Mech., 780 (2015), pp. 370–387]. Here, we rigorously establish the existence of FW-BICs in a broader class of cavity geometries by introducing perturbations to the refractive index under regularity constraints. Assuming the waveguide width is sufficiently small, we show that BICs correspond to intersections of two curves derived implicitly from the governing equations constructed via the mode-matching method. Crucially, we prove that such intersections are guaranteed, provided that two eigenvalues of the cavity intersect transversally and the associated eigenfunctions exhibit non-vanishing coupling to the radiation channel at the cavity-waveguide interface. Furthermore, our approach remains applicable for studying the emergence of FW-BICs under parameter-dependent boundary perturbations to the cavity.
keywords:
bound states in the continuum, Helmholtz equations, perturbation theory
{MSCcodes}
35Q60, 35P10, 78A45, 78M35
@@ 1 Introduction
We consider a broader class of 2D cavities, each with a semi-infinite waveguide attached to a segment of its boundary. The scattering matrix for wave systems in such structures was previously studied by Adamyan et. al. [1]. Let $h$ denote the waveguide width and $\mu$ the square of the free-space wavenumber. As $h\to 0$, the resonances $\lambda_{\mathrm{res}}$} of these open systems converge to the eigenvalues $\lambda$ of the corresponding closed cavities. The authors employ a two-step strategy, pioneered by H. Poincaré [21], to reveal the singular term $(\overline{{\lambda_{\mathrm{res}}}}-\mu)/({\lambda_{\mathrm{res}}}-\mu)$ in the scattering matrix.
In this work, we focus on establishing sufficient conditions for the existence of FW-BICs in the cavity-waveguide structures. Introducing an analytic perturbation to the refractive index parameterized by $\delta$, and assuming $h\ll 1$, we reduce the infinite-dimensional system derived from the mode-matching method to two governing equations for $\delta$ and $\mu$ in small intervals. Each equation defines an implicit function $\mu(\delta)$, and BICs are identified at the intersections of these curves in the $(\delta,\mu)$-plane. We demonstrate that such intersections are guaranteed to exist provided that two cavity eigenvalues intersect transversally as $\delta$ varies, and the associated eigenfunctions exhibit non-vanishing coupling to the radiation channel at the cavity-waveguide interface. The main result is formalized in Theorem 5.13. Our approach remains valid for demonstrating the existence of BICs under parameter-dependent boundary perturbations, if the eigenvalues and eigenfunctions exhibit continuous dependence on these perturbations. This continuity condition is satisfied in specific scenarios (cf. Section VII.6.5 in [11]), including scaling transformations of rectangular cavities as explored in [18]. For additional applications of analogous mode-matching techniques in resonance analysis of subwavelength structures, we refer the readers to [15, 17, 37, 36].
The structure of this paper is as follows. Section 2 is devoted to the problem formulation and the three key conditions underlying the analysis. The derivation of the governing infinite-dimensional linear system is presented in Section 3. Section 4 details the reformulation of this system into a real-valued form and its subsequent reduction to a finite-dimensional system. A rigorous proof for the existence of FW-BICs, which is the central result of this paper, is established in Section 5. Numerical simulations validating the theoretical findings are provided in Section 6. The paper concludes in Section 7 with a discussion on the extension of these results to boundary perturbations and other generalizations of the approach.
@@ 2 Problem formulation and key requirements
As illustrated in Fig. 1, our problem region $\Omega$ is defined as $\Omega:=\Omega_{\mathrm{in}}\cup\Gamma_{h}\cup\Omega_{\mathrm{out}}$, where $\Omega_{\mathrm{in}}$ represents the cavity region with a line segment on its boundary, $\Omega_{\mathrm{out}}$ denotes the waveguide region of width $h$ extended from the line segment, and $\Gamma_{h}$ is the common boundary between these two regions. Letting $\mathbf{o}$ denote the middle point of $\Gamma_{h}$, we establish a rectangular coordinate system in $\mathbb{R}^{2}$ centered at $\mathbf{o}$. Without loss of generality, we always fix the point $\mathbf{o}$ on $\partial\Omega_{\mathrm{in}}$ as the middle point of $\Gamma_{h}$ when $h$ is allowed to decrease in the subsequent sections. Additionally, $\Omega_{B}$ denotes a rectangular region in $\Omega_{\mathrm{in}}$ such that $\Gamma_{h}\subset\partial\Omega_{B}$.
[…]
@@
[…]
@@
(3) $-\mathcal{P}{u}+\mu{u}$ $=0\ \mathrm{in}\ \Omega,\$
@@
(5) $-\mathcal{P}{\psi}+\lambda{\psi}$ $=0\ \mathrm{in}\ \Omega_{\mathrm{in}},\$
@@
    At $\delta=0$, the eigenvalue functions $\lambda_{M-2}(\delta)$ and $\lambda_{M-1}(\delta)$ intersect transversally, resulting in an eigenvalue of multiplicity 2.
@@ 3.1 Eigenfunction expansion in the cavity
Under Condition A.1, we can choose a set of normalized eigenfunctions $\{\psi_{m}(\cdot,\delta)\}$ such that the mapping $\delta\to\psi_{m}(\cdot,\delta)$ is continuous from $\delta_{0},\delta_{0}]$ to $H^{1}(\Omega_{\mathrm{in}})$ for each $m\in\mathbb{N}$, i.e.,
(8) $\|\psi_{m}(\cdot,\delta_{a})-\psi_{m}(\cdot,\delta_{b})\|_{H^{1}(\Omega_{\mathrm{in}})}\to 0\ \text{as}\ \delta_{a}\to\delta_{b}\ \text{for}\ \delta_{a},\delta_{b}\in[-\delta_{0},\delta_{0}].$
@@
(9) $\left|\frac{1}{\lambda_{m}(\delta_{a})+1}-\frac{1}{\lambda_{m}(\delta_{b})+1}\right|\leq{e^{C_{0}|\delta_{a}-\delta_{b}|}-1},\ \text{for}\ \delta_{a},\delta_{b}\in[-\delta_{0},\delta_{0}],$
@@
Since $n^{2}(\cdot,\delta)$ varies analytically with respect to $\delta$ and $H^{1}(\Omega_{\mathrm{in}})$ is dense in $L^{2}(\Omega_{\mathrm{in}})$, $\Phi_{\delta}$ is a self-adjoint holomorphic family of forms of type (a) (cf. Sections VII.4.2 and VII.4.8 in [11]). Hence, $\psi_{m}(\cdot,\delta)$ can be chosen as a continuous mapping from $\delta_{0},\delta_{0}]$ to $L^{2}(\Omega_{\mathrm{in}})$, and its eigenvalue $\lambda_{m}(\delta)$ satisfies (9) by using the results in Section VII.4.7 of [11]. To prove (8), it suffices to show for each $m\in\mathbb{N}$ that
(12) $\left\|\frac{1}{n(\cdot,\delta_{b})}\nabla(\psi_{m}(\cdot,\delta_{a})-\psi_{m}(\cdot,\delta_{b}))\right\|_{[L^{2}(\Omega_{\mathrm{in}})]^{2}}\to 0\ \text{as}\ \delta_{a}\to\delta_{b}.$
A direct computation yields:
$\left\|\frac{1}{n(\cdot,\delta_{b})}\nabla(\psi_{m}(\cdot,\delta_{a})-\psi_{m}(\cdot,\delta_{b}))\right\|^{2}_{[L^{2}(\Omega_{\mathrm{in}})]^{2}}$
$=$ $(\frac{1}{n^{2}(\cdot,\delta_{a})}\nabla\psi_{m}(\cdot,\delta_{a}),\nabla(\psi_{m}(\cdot,\delta_{a})-\psi_{m}(\cdot,\delta_{b})))_{\Omega_{\mathrm{in}}}$
$-(\frac{1}{n^{2}(\cdot,\delta_{b})}\nabla\psi_{m}(\cdot,\delta_{b}),\nabla(\psi_{m}(\cdot,\delta_{a})-\psi_{m}(\cdot,\delta_{b})))_{\Omega_{\mathrm{in}}}$
$+((\frac{1}{n^{2}(\cdot,\delta_{b})}-\frac{1}{n^{2}(\cdot,\delta_{a})})\nabla\psi_{m}(\cdot,\delta_{a}),\nabla(\psi_{m}(\cdot,\delta_{a})-\psi_{m}(\cdot,\delta_{b})))_{\Omega_{\mathrm{in}}}$
$=$ $(\lambda_{m}(\delta_{a})\psi_{m}(\cdot,\delta_{a})-\lambda_{m}(\delta_{b})\psi_{m}(\cdot,\delta_{b}),\psi_{m}(\cdot,\delta_{a})-\psi_{m}(\cdot,\delta_{b}))_{\Omega_{\mathrm{in}}}$
(13) $+((\frac{1}{n^{2}(\cdot,\delta_{b})}-\frac{1}{n^{2}(\cdot,\delta_{a})})\nabla\psi_{m}(\cdot,\delta_{a}),\nabla(\psi_{m}(\cdot,\delta_{a})-\psi_{m}(\cdot,\delta_{b})))_{\Omega_{\mathrm{in}}}.$
Here, the second equality follows from the divergence theorem and the equations (5)–(6) satisfied by the eigenpair $(\lambda_{m},\psi_{m})$. Consequently, the right-hand-side of (13) converges to 0 as $\delta_{a}\to\delta_{b}$ by the $L^{2}$-convergence $\|\psi_{m}(\cdot,\delta_{a})-\psi_{m}(\cdot,\delta_{b})\|_{L^{2}(\Omega_{\mathrm{in}})}\to 0$ and the assumption $\|n(\cdot,\delta_{a})-n(\cdot,\delta_{b})\|_{L^{\infty}(\Omega_{\mathrm{in}})}\to 0$.
@@
(18) $\|u\|_{H^{1}(\Omega_{\mathrm{in}})}=\sqrt{\|u\|^{2}_{L^{2}(\Omega_{\mathrm{in}})}+\|\frac{1}{n}\nabla{u}\|^{2}_{[L^{2}(\Omega_{\mathrm{in}})]^{2}}},\ \text{for}\ u\in{H}^{1}(\Omega_{\mathrm{in}}).$
@@ 3.2 Mode-expansion in the waveguide
and $u$ is supposed to be outgoing as $x_{1}\to\infty$. This gives the expansions:
(21) $u=\sum_{j=0}^{\infty}b_{j}e^{{\bf i}\alpha_{j}x_{1}}\phi_{j}(x_{2})\ \text{in}\ H^{1}_{\mathrm{loc}}(\Omega_{\mathrm{out}}),$
@@
and $\alpha_{j}=\sqrt{\mu-\left(\frac{j\pi}{h}\right)^{2}}$. We also introduce a set of functions $\{\eta_{j}\}$
(23) $\eta_{j}:=\sqrt{\frac{2}{h}}\sin\left(\frac{j\pi}{h}x_{2}+\frac{j\pi}{2}\right)\ \text{for}\ j\in\mathbb{Z}_{+}.$
@@
(24) $u|_{\Gamma_{h}}=$ $\sum_{j=0}^{\infty}b_{j}\phi_{j}\$$\text{in}\ H^{1/2}(\Gamma_{h}),$
(25) $\partial_{\bm{\nu}}u|_{\Gamma_{h}}=$ $\sum_{j=0}^{\infty}b_{j}{\bf i}\alpha_{j}\phi_{j}\$$\text{in}\ \widetilde{H}^{-1/2}(\Gamma_{h}).$
@@ 4 Reduction of the system
In this section, we reduce the system (29) to a real-valued, finite-dimensional form by assuming $h$ is small enough and restricting $\delta$ and $\mu$ in small intervals.
@@
Let $M\geq 3$ be an integer and suppose the eigenvalues satisfy
(31) $\lambda_{M-3}(0)<\lambda_{M-2}(0)\ \text{and}\ \lambda_{M-1}(0)<\lambda_{M}(0).$
Then, under Condition A.1, for every $\epsilon>0$ there exists $\delta_{1}>0$ such that for all $\delta\in[-\delta_{1},\delta_{1}]$, the following bounds hold:
@@
By selecting $\delta_{1}$ such that
(36) $e^{C_{0}\delta_{1}}-1\leq\frac{1}{\lambda_{M}(0)+1-\epsilon}-\frac{1}{\lambda_{M}(0)+1},$
@@
(37) $\epsilon<\frac{1}{4}\min\Big\{\lambda_{M-2}(0)-\lambda_{M-3}(0),\lambda_{M}(0)-\lambda_{M-1}(0)\Big\}.$
We assume $\delta$ varies within a range $\delta_{2},\delta_{2}]$ for some $\delta_{2}>0$ such that (32)–(33) hold and
@@
For the subsequent analysis, we define the endpoints as $\mu_{l}:=\lambda_{M-3}(0)+2\epsilon$ and $\mu_{r}:=\lambda_{M}(0)-2\epsilon$. We also summarize the various parameter bounds used in the following sections as they are introduced for clarity. The intervals are:
  • $(0,h_{0})$: introduced in (48).
  • $(0,h_{1})$: introduced in Theorem 4.9.
  • $(0,h_{2})$ and $\delta_{3},\delta_{3}]$: introduced in Theorem 5.3.
  • $(0,h_{3})$: introduced in Theorem 5.11.
  • $(0,h_{4})$: introduced in Theorem 5.13.
Additionally, we require $h_{4}\leq{h_{3}}\leq{h_{2}}\leq{h_{1}}\leq{h_{0}}$ and $\delta_{3}\leq{\delta_{2}}\leq\delta_{1}\leq\delta_{0}$.